OpenStudy (anonymous):

Write the expression as a monomial containing a single fraction or a constant Sinx (cot^2x+1)

OpenStudy (raden):

hint : cot^2 (x) + 1 = csc^2 x

OpenStudy (raden):

so, sinx (cot^2x+1) = sinx * csc^2 x = sinx * 1/sin^2 x = .... (simplify)